# MKF-ADS: A Multi-Knowledge Fused Anomaly Detection System for Automotive

### 07 Mar 2024 v1

@@ Abstract
With the requirements of Intelligent Transport Systems (ITSs) for extensive connectivity of Electronic Control Units (ECUs) to the outside world, safety and security have become stringent problems. Intrusion detection systems (IDSs) are a crucial safety component in remediating Controller Area Network (CAN) bus vulnerabilities. However, supervised-based IDSs fail to identify complexity attacks and anomaly-based IDSs have higher false alarms owing to capability bottleneck. In this paper, we propose a novel multi-knowledge fused anomaly detection model, called MKF-IDS. Specifically, the method designs an integration framework, including spatial-temporal correlation with an attention mechanism (STcAM) module and patch sparse-transformer module (PatchST). The STcAM with fine-pruning uses one-dimensional convolution (Conv1D) to extract spatial features and subsequently utilizes the Bidirectional Long Short Term Memory (Bi-LSTM) to extract the temporal features, where the attention mechanism will focus on the important time steps. Meanwhile, the PatchST captures the combined long-time historical features from independent univariate time series. Finally, the proposed method is based on knowledge distillation to STcAM as a student model for learning intrinsic knowledge and cross the ability to mimic PatchST. In the detection phase, the MKF-ADS only deploys STcAM to maintain efficiency in a resource-limited IVN environment. Moreover, the redundant noisy signal is reduced with bit flip rate and boundary decision estimation. We conduct extensive experiments on six simulation attack scenarios across various CAN IDs and time steps, and two real attack scenarios, which present a competitive prediction and detection performance. Compared with the baseline in the same paradigm, the error rate and FAR are 2.62% and 2.41% and achieve a promising F1-score of 97.3%.
@@
Intelligent Transportation System, Control Area Network, In-Vehicle Security, Anomaly Detection, Knowledge Fused
@@ I Introduction
Compared with traditional vehicles, intelligent connected vehicles (ICVs) have been an enormous advance, motivating intelligent transport systems (ITSs) to develop a significant step [1]. The fundamental purpose of ITS is to make driving tasks easier by minimizing the demands on the driver [2]. In order to achieve such goals, various services related to vehicle intelligence are being extensively researched, such as autonomous vehicles (AVs), vehicular network communication, and cloud computing [3, 4]. However, these intelligent services also make the otherwise closed communication in-vehicle networks (IVNs) confront more attacks from the inter-extra [5, 6]. These attacks could have serious repercussions since vehicles maliciously manipulated will pose a serious risk to the lives of drivers and pedestrians [7, 8]. Therefore, vehicular network security is also a crucial research area.
@@
The modern automotive communication protocol standard is the bus-based control area network (CAN) because of its advantages of low cost, minimal latency, and high efficiency [9]. Unfortunately, the CAN bus was not designed with any security mechanisms in mind (e.g., authentication and encryption) [10], which makes a large number of electronic control units (ECUs) equipped in vehicles vulnerable to security attacks when communicating via the bus or with external devices via Bluetooth or hardware ports [3, 11]. This is the fundamental reason why vehicle cyber security must be considered before implementing the aforementioned intelligent services. While a symptomatic solution to security problems (e.g., implementing encryption or authentication) is the most straightforward, more work has proven that such approaches consume a lot of memory and computation [12, 13, 14]. Moreover, such approaches require consideration of backward compatibility concerns [15]. In contrast, the in-vehicle intrusion detection systems (IV-IDS) are an efficient protection mechanism used to detect cyber-attacks on automobiles [16], which has two main categories, signature-based and anomaly-based [17, 18, 15]. Signature-based IDS use supervised learning to capture fixed types of attacks from existing databases. However, the method requires a large amount of marker data and real-life CAN attack data is not readily available. As anomaly-based IDS solutions can detect unknown malicious network activities, especially zero-day attacks, it is more suitable and efficient for the complex and open network environment of modern vehicles [19].
@@
Some ADS research has begun to model the inherent correlation of the normal CAN messages, so as to identify a deviation in detecting malicious messages. For instance, Song et al. [20] presented an anomaly method that identifies attack messages through noisy pseudo-normal data. However, this method is supervised and thus requires abnormal datasets in the learning process. In contrast, Talyor et al  and Qin et al. are both used for long and short-term memory (LSTM) to learn the temporal correlation of CAN transition bits, in order to search for a deviation between abnormal and normal messages. But they made the model at all bits (usually 64 bits) without considering the actual meaning, resulting in the predicted performance being relatively worse. Meanwhile, the Convolutional Neural Network (CNN) and attention mechanism have been introduced as a state-of-the-art paradigm aimed at capturing comprehensive features [21, 22]. However, they fail to capture long time-series historical dependencies. Recently, Alkhatib et al. [23] designed a CAN-BERT model that learns the context semantic features of CAN ID sequences. Nam et al. [24] regarded the CAN ID sequence as a sentence with several words and adopted the Generative Pretrained Transformer (GPT) model to learn the normal pattern of each sequence. Although contextual features are important for ADS, these advanced transformer-based models with their large number of parameters are not able to satisfy IVNs with limited resources.
@@
To address the above challenges, we propose a Multi-Knowledge Fused Anomaly Detection System, called MKF-ADS. The method detects malicious attacks from deviations learned from normal messages, which follows the paradigm of multivariate time series prediction. To build a more actual meaningful model, we used the “READ” method to preprocess the CAN payload. This method extracts the signal boundary and continuous physical signal to assist our model drop some noisy bits. MKF-IDS is unsupervised to capture valuable features on normal CAN data that is a multivariate time series and predict the message on the next timestamp. The deviation is built on the actual signals and prediction signals. To capture spatial-temporal correlation features, the MKF-ADS fuse one-dimensional CNN and bidirectional LSTM (Bi-LSTM) with a soft attention mechanism (STcAM), where the parameters are much less than the previous methods. To retain local semantic information and enhance the long-time-series historical dependency, we introduce a patch sparse-transformer network (PatchST) to identify context features from separated single time series. Considering the resource limit of IVN, we fused spatial-temporal correlation features and context features based on a cross-knowledge distillation, where the STcAM as student model absorbs knowledge from PatchST as the teacher model. Thus, the final MKF-IDS is lightweight and contains only the STcAM module with comprehensive knowledge inference capability. Extensive experiments have shown that the proposed model can reduce the error rate and FAR to 2.62% and 2.41%, respectively. Compared with the baseline as the same paradigm, our method achieves a competitive F1-score of 97.3%.
The main contributions of this paper are as follows.
@@
    We propose a novel neural network CAN anomaly prediction model MKF-ADS by fusing multi-knowledge, of which spatial-temporal correlation important features are extracted by the STcAM module, and long time-series historical dependency knowledge is absorbed from the transformer-based module PatchST.
@@
    We construct a multi-knowledge fused framework, which regards the StcAM as a student model and the PatchST as a teacher model. STcAM not only learns inherent knowledge but also cross-mimicks valuable features from PathchST, which avoids conflict on different knowledge and improves the training stability.
@@
    We evaluate the prediction and detection performance of six intrusion scenarios and then generalization performance is verified on different CAN IDs, multi-time steps, and real-attack tasks in ROAD tasks. Also, The feasibility of the model’s complexity and overall efficiency is discussed when applied in the automotive.
@@ II Related Work
Technology Entropy and frequency HMM, SVM, and ARIMA CNN, LSTM, GRU, and AM CNN, LSTM, and One-class CNN, LSTM, and AM BERT and GPT Multi-knowledge fused
@@ II-A Security Models for Vehicular Network
Authentication-based and cryptography-based methods have been studied over the years by several researchers [42, 43]. Nilsson et al. first used authentication mechanisms to prevent IVN attacks, which must insert 64-bit message authentication codes (MACs) by a 16-bit tag consisting of four messages in the CRC field [44]. Thereafter, an out-of-band channel is presented to transmit authentication messages in [45], and variable-length MACs are proposed in [46]. In short, they are either computationally demanding or unsafe, which prevents them from providing dependable frame-by-frame authentication. Han et al. presented to divide the ECU’s functional attributes by the features of ICV, and then realize security isolation communication via the one-time-pad algorithm [47]. Moreover, a combination method of encryption and authentication mechanisms was provided in [48], and in [49] uses a dynamic symmetric key to encrypt the 8byte data payload but does not modify the CAN ID. Although these methods also provide a certain security guarantee for IVN, they also suffer from response delays or increased bus loads [50]. Hence, providing security defense via MACs or encryption is not a straightforward proposition.
@@
Intrusion Detection Systems (IDSs) have been widely studied owing to taking appropriate countermeasures in the event of attacks. Early IVN-IDSs identify malicious messages through a predefined knowledge base (e.g., frequency and entropy) [25, 26, 27, 28], but these methods have a high false positive rate (FPR) and cannot detect unknown attacks. Generally, statistics-based IDS are efficient and lightweight but may ignore low-volume and aperiodic frequency attacks. Machine learning-based methods have also been extensively introduced for intrusion detection. SVM [29], HMM [30], and ARIMA [31] are researched to protect IVN security. However, these approaches have limitations in detection performance. Due to the great density of the generated vehicular data and enhanced computing resources, various deep learning-based strategies have recently been presented that outperform the current signature-based or machine-learning approaches. Song et al[32] designed a deep convolution neural network (DCNN) based method for anomaly detection, removing the need for extracting handcrafted features. To extract valuable features, the CNN and LSTM-based deep learning model was presented by Xun et al[33]. The hybrid model was also introduced in  [34]. Further, Cheng et al[35] incorporated attention mechanisms on this basis to reduce the false alarm rate. Javed et al[36] replaced the time-permissive feature extraction model LSTM with the GRU model to improve the inference speed. Although the detection capabilities of these methods have improved to varying degrees, they all rely heavily on a large number of pre-labeled labels for supervised learning and are not practical in defending against attack variants.
@@
To this end, anomaly detection systems (ADSs) have been struggling to break through on how to model unknown attacks. Ashraf et al[7] utilized auto-encoder architecture to reconstruct the input messages via spatial-temporal features. Similarly, Agrawal et al[3] presented a method that could reconstruct directly using spatial-temporal features and then detect anomalies based on threshold deviations, which are simpler and more intuitive than existing auto-encoder architectures. Tariq et al[37] trained a Convolutional LSTM-based model by transfer learning to identify attacks, called CANTrasfer. They also applied one-shot learning so the system could detect intrusion attacks adaptively variable datasets, but the detection performance is relatively poor. Zhang et al. [38] presented graph convolutional neural network (GCNN)-based ADS to capture the falsification attacks. Considering that CAN messages always follow comparatively fixed patterns, ADS can use predictive analysis to detect complex cyber attacks. This paradigm is stringent excluding all suspicious messages that deviate from the prediction. Taylor et al. [39] and Qin et al. [40] developed the LSTM-based model to predict the next package, where the error measure is made using the predicted output value of the log loss function. Sun et al. [22] introduced convolution and attention mechanism (AM) to make up for valuable feature extraction. Interestingly, In  [23] and  [24], the outstanding model in NLP has been suggested applying to ADS (e.g. Bidirectional Encoder Representations from Transformers (BERT) or Generative pre-trained Transformer (GPT)). Although the plentiful context features are incorporated, the number of model parameters is tens to hundreds of times higher than the previous scheme.
@@ II-B Comparation In-Depth for IVN-IDS
To further present our advantages, the main characteristics and functionalities of the MKF-ADS are listed in Table  I. Table  I also compares existing works with the proposed model based on the significant requirements for IVN-IDS.
According to the modeling strategies and objectives, these works are divided into Statistical Learning (SL), Machine Learning (ML), and Deep Learning (DL). The majority of studies focus on DL-based and build different objectives, including supervised learning to identify known attacks, and anomaly-based to classify or predict unknown attacks. Intuitively, we find that early research based on SL and ML have similar characteristics that fixed feature extraction patterns  [27, 25, 28] and worrisome performance [30, 29]. In contrast, supervised-based DL is substantially improved by rich feature extraction. For example, CNN extracts spatial features [32], LSTM extracts temporal features [34], and the attention mechanism models valuable features [35]. Despite this, the shortcomings are evident, e.g., higher computation complexity and limited detection range. When the attention is transferred to anomaly classification, reconstruction modeling aims to attain the deviation score of normal messages [3, 7]. In general, this technique suffers from significant collection delays. Also, the one-class classifier is a bare-bones migration that always sacrifices the detection ability of known attacks [37, 38].
@@
Taking into account the complementarity of spatial-temporal and contextual features, the multi-knowledge fused is practical. Most importantly, for deploying IDS in real vehicle systems, lower computation complexity is required to satisfy real-time detection. However, only a limited amount of work has been performed on workstations for simulation testing [27, 32]. Some works [9, 28, 35] have performed vehicle-level testing or real-time analysis. Hence, our proposed MKF-ADS constructs a cross-knowledge distillation framework that regards contextual knowledge as a teacher and spatial-temporal knowledge as a student. We hope that the proposed model can capture subtle deviations through enriching knowledge and then identify potential threats. Notably, we introduce sharing components to capture spatial-temporal knowledge, and the reduced capabilities are instructed by contextual knowledge, which is a guarantee for real-time detection. Also, robustness validation is presented by fine-tuning the proposed MKF-ADS.
@@ III-A Control Area Network Bus
The Control Area Network (CAN) bus is a de facto communication standard for vehicular networks, which divides ECUs into different domains, e.g., safety domain, power train domain, and comfort domain. Each of these communication sub-networks is responsible for different key tasks such as automatic crash notification, vehicle diagnostics, and infotainment services  [51]. To achieve intelligent services, more external interfaces (e.g., communication via WIFI, Bluetooth, and 5G) are opened up, which also increases the potential vulnerability exposure.

### 15 Mar 2024 v2

# MKF-ADS: Multi-Knowledge Fusion Based Self-supervised Anomaly Detection System for Control Area Network

@@ Abstract
Control Area Network (CAN) is an essential communication protocol that interacts between Electronic Control Units (ECUs) in the vehicular network. However, CAN is facing stringent security challenges due to innate security risks. Intrusion detection systems (IDSs) are a crucial safety component in remediating Vehicular Electronics and Systems vulnerabilities. However, existing IDSs fail to identify complexity attacks and have higher false alarms owing to capability bottleneck. In this paper, we propose a self-supervised multi-knowledge fused anomaly detection model, called MKF-ADS. Specifically, the method designs an integration framework, including spatial-temporal correlation with an attention mechanism (STcAM) module and patch sparse-transformer module (PatchST). The STcAM with fine-pruning uses one-dimensional convolution (Conv1D) to extract spatial features and subsequently utilizes the Bidirectional Long Short Term Memory (Bi-LSTM) to extract the temporal features, where the attention mechanism will focus on the important time steps. Meanwhile, the PatchST captures the combined contextual features from independent univariate time series. Finally, the proposed method is based on knowledge distillation to STcAM as a student model for learning intrinsic knowledge and cross the ability to mimic PatchST. We conduct extensive experiments on six simulation attack scenarios across various CAN IDs and time steps, and two real attack scenarios, which present a competitive prediction and detection performance. Compared with the baseline in the same paradigm, the error rate and FAR are 2.62% and 2.41% and achieve a promising F1-score of 97.3%.
@@
Control Area Network, Anomaly Detection System, Spatial-temporal features, Contextual features, Knowledge distillation.
@@ I Introduction
Control Area Network (CAN) has become the standard communication protocol for vehicle networks, playing a decision-making role in the interaction of multiple Electronic Control Units (ECUs) [1]. The emergence of Intelligent Connected Vehicles (ICVs) has led researchers to focus more on CAN. Therein, interconnectivity requires CAN to open additional interfaces to achieve intelligent services, such as autonomous vehicles (AVs), vehicular network communication, and cloud computing [2, 3, 4]. However, these intelligent services also make the otherwise closed communication in-vehicle networks (IVNs) confront attacks from the inter-extra [5]. Unfortunately, the CAN bus was not designed with any security mechanisms in mind (e.g., authentication and encryption) [6]. When communicating via the bus or with external devices via Bluetooth or hardware ports [7, 8, 9], the bus or ECUs may be hijacked, which will pose serious repercussions to the lives of drivers and pedestrians [10, 11, 12]. Therefore, vehicular network security is a crucial research area.
@@
Although taking purpose-direction to security problems is the most straightforward (e.g., encryption or authentication) [13], more work has proven that such approaches consume a lot of memory and computation [14, 15, 16]. Moreover, such approaches require consideration of backward compatibility concerns [17]. In contrast, the in-vehicle intrusion detection systems (IV-IDS) are an efficient protection mechanism used to detect cyber-attacks on automobiles [18], which has two main categories, signature-based and anomaly-based [19, 20]. Signature-based IDS use supervised learning to capture fixed types of attacks from existing databases. However, the method requires a large amount of marker data, while real-life CAN attack data is not readily available. As anomaly-based IDS can detect unknown malicious network activities, especially zero-day attacks, it is more suitable for the open and complex vehicular network environment [21].
@@
Some ADS research has begun to model the inherent correlation of the normal CAN messages, so as to identify a deviation in detecting malicious messages. For instance, Song et al. [22] presented an anomaly method that identifies attack messages through noisy pseudo-normal data. However, this method is supervised and thus requires abnormal datasets in the learning process. In contrast, Talyor et al [23] and Qin et al. are both used for long and short-term memory (LSTM) to learn the temporal correlation of CAN transition bits, in order to search for a deviation between abnormal and normal messages. But they made the model at all bits (usually 64 bits) without considering the actual meaning, resulting in the predicted performance being relatively worse. Meanwhile, the Convolutional Neural Network (CNN) and attention mechanism have been introduced as a state-of-the-art paradigm aimed at capturing comprehensive features [24, 25]. However, they fail to capture long time-series historical dependencies. Recently, Alkhatib et al. [26] designed a CAN-BERT model that learns the context semantic features of CAN ID sequences. Nam et al. [27] regarded the CAN ID sequence as a sentence with several words and adopted the Generative Pretrained Transformer (GPT) model to learn the normal pattern of each sequence. Although contextual features are important for ADS, these advanced transformer-based models cannot satisfy IVNs with limited resources due to the large number of parameters.
@@
To address the above challenges, we propose a Multi-Knowledge Fusion based self-supervised Anomaly Detection System, called MKF-ADS. The method is self-supervised to detect malicious attacks from deviations learned from normal messages, which follows the paradigm of multivariate time series prediction. The deviation is built on the actual signals and prediction signals. In order to construct a more realistic and meaningful model, the CAN payload is preprocessed using the ”READ” method, thereby determining the signal boundaries and extracting the continuous physical signal. The decision reference of MKF-ADS is from multi-knowledge, consisting of spatial-temporal correlation and contextual information. The fusion mechanism is a guidance and imitation mode from high to low due to the resource limit of IVN. Extensive experiments have shown that the proposed model can reduce the error rate and FAR to 2.62% and 2.41%, respectively. Compared with the baseline as the same paradigm, our method achieves a competitive F1-score of 97.3%. The main contributions of this paper are as follows.
@@
    We propose MKF-ADS, a self-supervised anomaly prediction model by fusing multi-knowledge for CAN. The framework consists of the STcAM and PatchST modules, used for extracting spatial-temporal correlation knowledge and long time-series historical dependency knowledge.
@@
    We construct a multi-knowledge fusion framework, which regards the StcAM as a student model and the PatchST as a teacher model. STcAM not only learns inherent knowledge but also cross-mimicks valuable features from PathchST, which avoids conflict on different knowledge and improves the training stability.
@@
    We evaluate the prediction and detection performance of six intrusion scenarios and then generalization performance is verified on different CAN IDs, multi-time steps, and real-attack tasks in ROAD tasks. Also, The feasibility of the model’s complexity and overall efficiency is discussed when applied in the vehicular network.
@@ II Related Work
Technology Entropy and frequency HMM, SVM, and XGBoost CNN, LSTM, GRU, and AM CNN, LSTM, and One-class CNN, LSTM, and AM BERT and GPT Multi-knowledge fused
@@ II-A Security Models for Vehicular Network
Authentication-based and cryptography-based methods have been studied over the years by several researchers [44, 45]. dariz et al. [46] provide a combined encryption and authentication mechanisms for vehicular network, and in [47] uses a dynamic symmetric key to encrypt the 8byte data payload but does not modify the CAN ID. Although these methods provide a certain security guarantee for IVN, they also suffer from response delays or increased bus loads [48]. Alladi et al. [49] proposed a lightweight and secure authentication and attestation scheme for attesting vehicles while they are on the roads. Cui et al. [50] proposed a lightweight encryption and authentication scheme for the CAN bus of AVs using message authentication codes and Grain stream cipher. In short, they are either computationally demanding or protocol-modified, which prevents them from providing dependable efficient defense. Hence, providing security defense via MACs or encryption is not a straightforward proposition.
@@
Intrusion Detection Systems (IDSs) have been widely studied owing to taking positive countermeasures in the event of attacks [51]. Early IVN-IDSs identify malicious messages through a predefined knowledge base (e.g., frequency and entropy) [29, 30, 31], but these methods have a high false positive rate (FPR) and cannot detect unknown attacks. Generally, statistics-based IDS are efficient and lightweight but may ignore low-volume and aperiodic frequency attacks. Machine learning-based methods have been extensively researched [32, 33]. However, these approaches have a botnack in detection performance. Due to the great density of the generated vehicular data and enhanced computing resources, various deep learning-based strategies have recently been presented that outperform the current signature-based or machine-learning approaches. Song et al[34] designed a deep convolution neural network (DCNN) based method for anomaly detection, removing the need for extracting handcrafted features. To extract valuable features, the CNN and LSTM-based deep learning model was presented by Xun et al[35]. The hybrid model was also introduced in  [36]. Further, Cheng et al[37] incorporated attention mechanisms on this basis to reduce the false alarm rate. Javed et al[38] replaced the time-permissive feature extraction model LSTM with the GRU model to improve the inference speed. Although the detection capabilities of these methods have improved to varying degrees, they all rely heavily on a large number of pre-labeled labels for supervised learning and are not practical in defending against attack variants.
@@
Anomaly detection systems (ADSs) have been struggling to break through on how to model unknown attacks. Ashraf et al[10] utilized auto-encoder (AE) to reconstruct the input messages via spatial-temporal features. Similarly, Agrawal et al[2] presented a method that reconstructs sequence directly using spatial-temporal features and then detects anomalies based on threshold deviations. Jeong et al. [39] present explainable intrusion detection based on the self-supervised AE paradigm. Tariq et al[40] trained a Convolutional LSTM-based model by transfer learning to identify attacks, called CANTrasfer. They also applied one-shot learning so the system could detect adaptive attacks, but the detection performance is relatively poor. Zhang et al. [41] presented graph convolutional neural network (GCNN)-based ADS to capture the falsification attacks. Considering that CAN messages always follow comparatively fixed patterns, ADS can use predictive analysis to detect complex cyber attacks. This paradigm is stringent excluding all suspicious messages that deviate from the prediction. Taylor et al. [23] and Qin et al. [42] developed the LSTM-based model to predict the next package, where the error measure is made using the predicted output value of the log loss function. Sun et al. [25] introduced convolution and attention mechanism (AM) to make up for valuable feature extraction. Interestingly, In  [26] and  [27], the outstanding model in NLP has been suggested applying to ADS (e.g. BERT [26] and GPT [27]). Although the plentiful context features are incorporated, the number of model parameters is tens to hundreds of times higher than the previous scheme.
@@ II-B Comparation In-Depth for IVN-IDS
To further present our advantages, the main characteristics and functionalities of the MKF-ADS are listed in Table  I. According to the modeling strategies and objectives, these works are divided into Statistical Learning (SL), Machine Learning (ML), and Deep Learning (DL). The majority of studies focus on DL-based and build different objectives, including supervised learning to identify known attacks, and anomaly-based to classify or predict unknown attacks. Intuitively, we find that early research based on SL and ML have similar characteristics that fixed feature extraction patterns  [28, 30, 31] and worrisome performance [32]. In contrast, supervised-based DL is substantially improved by rich feature extraction. For example, CNN extracts spatial features [34], LSTM extracts temporal features [36], and the attention mechanism models valuable features [37]. Despite this, the shortcomings are evident, e.g., higher computation complexity and limited detection range. When the attention is transferred to anomaly classification, reconstruction modeling aims to attain the deviation score of normal messages [2, 10]. In general, this technique suffers from significant collection delays. Also, the one-class classifier is a bare-bones migration that always sacrifices the detection ability of known attacks [40, 41].
@@
Taking into account the complementarity of spatial-temporal and contextual features, the multi-knowledge fused is practical. Most importantly, for deploying IDS in real vehicle systems, lower computation complexity is required to satisfy real-time detection. However, only a limited amount of work has been performed on workstations for simulation testing [30, 34]. Some works [7, 31, 37] have performed vehicle-level testing or real-time analysis. Hence, our proposed MKF-ADS constructs a cross-knowledge distillation framework that regards contextual knowledge as a teacher and spatial-temporal knowledge as a student. We hope that the proposed model can capture subtle deviations through enriching knowledge and then identify potential threats. Notably, we introduce fine-pruning to modeling spatial-temporal knowledge, and the reduced capabilities are instructed by contextual knowledge, which is a guarantee for real-time detection. Also, robustness validation is presented by fine-tuning the proposed MKF-ADS.
@@ III-A Control Area Network Bus
Control Area Network (CAN) bus is a de facto communication standard for vehicular networks, which divides ECUs into different domains, e.g., safety domain, power train domain, and comfort domain. Each of these communication sub-networks is responsible for different key tasks such as automatic crash notification, vehicle diagnostics, and infotainment services  [52]. To achieve intelligent services, more external interfaces (e.g., communication via WIFI, Bluetooth, and 5G) are opened up, which also increases the potential vulnerability exposure.